OpenStudy (anonymous):

@iPwnBunnies 5/8% as a fraction and a decimal in simplelist form…Sorry dyslexia

OpenStudy (ipwnbunnies):

Do you mean 5/8 as a decimal? Or really 5/8 percent?

OpenStudy (anonymous):

5/8 percent…That is what confuzzled me

OpenStudy (ipwnbunnies):

Ahh ok. Do you know what 5/8 is in decimal form?

OpenStudy (anonymous):

.625?

OpenStudy (ipwnbunnies):

Good. Now, when converting percentages to decimals, you want to move the decimal two places to the left.

OpenStudy (ipwnbunnies):

For example: If we have 50%, it'll be 0.5 in decimal form. Since we have 0.625%, what would that be in decimal form?

OpenStudy (anonymous):

.00625 I think

OpenStudy (ipwnbunnies):

Yes, 5/8% written as a decimal is 0.00625.

OpenStudy (anonymous):

62.5?

OpenStudy (ipwnbunnies):

Now, as a fraction. "percent" means out of 100

OpenStudy (ipwnbunnies):

For example: If we want to write 75% as a fraction, it would 75/100 = 3/4.

OpenStudy (anonymous):

thx SO MUCH! This has helped me because my teacher doesn't really explain things clearly…

OpenStudy (ipwnbunnies):

75 percent = 75 out of 100 parts.

OpenStudy (ipwnbunnies):

What did you get?

OpenStudy (anonymous):

.00625 AND 1/32

OpenStudy (ipwnbunnies):

That's not the correct fraction, try again. \[\frac{5}{8} \div 100 = \frac{5}{8}*\frac{1}{100}\]

OpenStudy (anonymous):

1/160? This is so complicated thx for helping. :)

OpenStudy (ipwnbunnies):

Correct, good job.

OpenStudy (anonymous):

thank you…I might have some more questions later… *hugz*

OpenStudy (anonymous):

write each fraction or decimal as a percent. Round to the nearest tenth of a percent if necessary… 7/500 AND 0.8/100… I'm really sorry- our teacher just doesn't really explain well

OpenStudy (ipwnbunnies):

This is easy. Find the decimal number of 7/500. Move the decimal to the right two places and add a percent sign. For the second fraction, remember I said percent is the same as out of 100? Since 0.8 is out of 100, it'll simply be...

OpenStudy (anonymous):

.008?

OpenStudy (ipwnbunnies):

Nooooo, remember: 75/100 = 75%; 342/100 = 342%. What is 0.8/100 in percentage?

OpenStudy (anonymous):

80… sorry

OpenStudy (ipwnbunnies):

What did you get for the first one, too.

OpenStudy (anonymous):

UR CONFUSING ME!

OpenStudy (ipwnbunnies):

Why would 0.8/100 = 80%? Again, percent is out of 100. So 0.8/100 = 0.8%

OpenStudy (anonymous):

FACEPALMS…lol gotcha